OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1

OpenStudy (anonymous):

y = 9

OpenStudy (anonymous):

you want me to show you how i did it?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok First you have to make the equation into like this 20x + 16y = 4 20x + 15y = -5 then you get the equation because you are lloking for the y you have to take out the x so 20x + 16y = 4 -20 - 15y = 5 so you get y = 9

OpenStudy (anonymous):

thank you i think i got it now.

OpenStudy (anonymous):

yw